3y^2/3+21y^1/3+30=0 Whats the solution to this? an can you show step by step so i'll know how to solve next time?
Answer by stanbon(75887)

3y^2/3+21y^1/3+30=0
---
Quadratic Form with variable y^(1/3).
---
Let w = y^(1/3) to get:
----
3w^2 + 21w + 30 = 0
3[w^2+7w+10] = 0
(w+5)(w+2) = 0
w = -5 or w = -2
-----
Now solve for y:
y^(1/3) = -5 or y^(1/3) = -2
y = -125 or y = -8
=====================
Cheers,
Stan H.